Diffusion Morning Report 05/13/09

Ted Berg tberg at stanford.edu
Wed May 13 06:29:06 PDT 2009

Good morning All,
    Looking good this morning. Ray has replaced the control vlave for H2 
on Tylan 3 and it needs testing. Tylan LTO temp is a bit off at 300 but 
good at 350. Thermco LTO testing. No other issues at this time.
  	Date: 	5/13/09 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	poor uniformity 	  	  	 
Tylan 3 	down 	needs test after repair 	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	up 	  	3.8/30 	  	 
Tylan 8-SiGe 	up 	  	15.54/30 	  	 
Tylan 9-Poly 	up 	  	20.77/30 	  	 
Tylan 10-Nitride 	up 	  	15.54/30 	  	 
Tylan 11-LTO 	yellow 	300 degrees is abit off 	2.4/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	up 	  	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco bpsg 	future 	testing for qual 	  	  	 
TEL Spa 	yellow 	  	  	  	 
Burn Box 	up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	A on line 548 psi 	 changeover working well  	  	 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20090513/1b6cbaeb/attachment.html>

More information about the diffusion mailing list